# Periodic Chains Scheduling on Dedicated Resources - A Crucial Problem in Time-Sensitive Networks

### v1

@@ 3.1.1 Single Period
Assume all chains have the same period $T$. Then, finding a feasible schedule for a single resource becomes an easy problem. We can prove that as for the two-machine flow-shop problem, minimizing the degeneracy can be done in polynomial time. Notice that in this case, the sum of processing times of tasks on any resource is not greater than $T$. Otherwise, the problem would be infeasible.
@@
Scattering on a line of two resources for $PD2|T_{i}=T,jit=0,chains|D_{sum}$ and $PD2|T_{i}=T,jit=0,chains|D_{\max}$ can be done in polynomial time.
@@
The non-periodic version of this problem is a two-machine flow-shop. A chain $\mathcal{C}^{k}$ corresponds to a job with two operations $J_{1}^{k},J_{2}^{k}$ performed on the first and second machine, respectively. And we know from [35], a schedule that optimizes the makespan ($C_{max}=\max_{k\in\{1,\ldots,\kappa\}}t_{2}^{k}+p_{2}^{k}$) can be computed in polynomial time using Johnson’s rule. Moreover, we know that it is a permutation schedule: the order of jobs is the same on both machines even though it was not required. If we shift the schedule of the first machine to the left and of the second machine to the right so that there does not remain idle time between two tasks on the resources and it ends at the optimal makespan, then the schedule can be repeated every $T$ time units without conflicts. We prove that this minimizes degeneracy.
Indeed, consider a pivot chain $\mathcal{C}^{k^{*}}$, for which $t_{1}^{k^{*}}+p_{1}^{k^{*}}=t_{2}^{k^{*}}$. Such a pivot chain exists. Otherwise, the schedule on the second machine could be left shifted, and thus, the makespan would not be optimal. The start-to-end latency for chain $\mathcal{C}^{k^{*}}$ is thus minimal, and so is its degeneracy. Then for any chain $k$ performed before this $k^{*}$ in the schedule, necessarily the start-to-end latency $S^{k}\leq t_{2}^{k^{*}}$ which is not greater than the sum of processing times on the first resource, which is not greater than $T$. Hence $D^{k}=0$. Symmetrically, for a chain $l$ performed after chain $k^{*}$ the value of $S^{l}$ should be less than the sum of processing times on the second resource, which is not greater than $T$ so that $D^{l}=0$. This proves that this schedule minimizes both $D_{\max},D_{sum}$. ∎
3.1.2 Two Periods
As in Section 3.1.1, we assume scattering chains on two resources, but now the chains can have two different harmonic periods. In [6], the authors give a reduction from $3-PARTITION$ to a single resource scheduling problem with two different harmonic periods, thus proving that the feasibility of a single resource PSP with two periods is strongly NP-Hard.
From this result, we can deduce the following:
Property 1.
For scattering chains on two resources in line topology, finding a schedule with $D_{sum}$ or $D_{\max}$ less than a value $d$ is strongly NP-hard.
@@ 3.1.2 Two Periods
Consider an instance of the single resource scheduling problem with two harmonic periods. Period $T^{i}$ of task $i$ is equal to one of the harmonic periods $\tau(1)$ or $\tau(2)$. Notice that the processing time of $i$ can be assumed strictly less than its period; otherwise $i$ would be the only task scheduled on the resource. For each task $i$ of the single resource problem, we define a chain $\mathcal{C}^{i}$ with period $T^{i}$ composed of task $i$ on the first resource and a task with processing time $1$ on the second resource. Obviously, if a feasible schedule for the single resource instance exists, it gives a schedule for the two resources of the chains with 00 degeneracy by scheduling the tasks on the second resource as soon as possible with respect to precedence constraint. Conversely, a schedule with 00 degeneracy implies the existence of a feasible schedule on the first resource. So this defines a reduction from the single resource, two harmonic periods problem to the scattering problem on two resources with two periods. ∎
As was shown in [6], the single resource problem with processing times and the periods belonging to a harmonic set is polynomial. However, the complexity of the two-resource scattering on a line case where periods and processing times are from one harmonic set remains open.
3.2 Scattering on a Line or an Out-Tree
In this section, we assume equal processing times of all tasks of a chain, and that the resources are organized along a given topology. We first consider the special case of scattering along a line.
3.2.1 Scattering on a Line
In this subsection, we consider a line of resources indexed from $1$ to $M$, and we suppose that each chain $\mathcal{C}^{k}$ starts at resource $1$ and ends at resource $u^{k}$. The chain comprises of tasks $J_{i}^{k}$ with $1\leq i\leq u^{k}$. Let us denote by $p^{k}$ the processing time of every task of chain $\mathcal{C}^{k}$.
So resource $1$ processes all the tasks from all chains $\mathcal{C}^{k},k=1,\ldots,\kappa$. Then, if there exists a feasible schedule for resource $1$, there exists a feasible schedule for each resource and thus a feasible schedule for the problem.
We call an offset schedule a periodic schedule built by assuming that for any resource $i$ a chain $\mathcal{C}^{k}$, task $J_{i}^{k}$ with $i\leq u^{k}$ is scheduled exactly as it is in resource $1$, with an offset that only depends on $i$. We define $o_{i}$ to be the offset between schedules of resource $1$ and resource $i$. Then, for each chain $\mathcal{C}^{k}$ with $u^{k}\geq i>1$ it holds that $t_{i}^{k}=t_{1}^{k}+o_{i}$.
Property 2.
Suppose we are given a feasible schedule $\sigma$ of the tasks on the first resource. The minimal degeneracy $D_{\max}$ or $D_{sum}$ of an offset schedule $\omega$ associated with $\sigma$ can be computed in polynomial time. In this schedule, we have the following inequality:
@@ 3.2.1 Scattering on a Line
Let $i>1$ be a resource. We can prove that in any feasible offset schedule, $o_{i}-o_{i-1}\geq\max_{k,u^{k}\geq i}p^{k}$. Thus $o_{i}\geq i\cdot\max_{k,u^{k}\geq i}p^{k}$. Indeed, let $\mathcal{C}^{k^{*}}$ be the chain with maximum processing time among chains with $u^{k}\geq i$. We know that $J_{i-1}^{k^{*}}$ precedes $J_{i}^{k^{*}}$ so that $t_{i}^{k}\geq t_{i-1}^{k}+p^{k^{*}}$. Notice that in any schedule, the start-to-end latency of a chain $\mathcal{C}^{k}$ equals $S^{k}=o_{u^{k}}+p^{k}$, So in any offset schedule, the minimum start-to-end latency of the chains is obtained by stating $o_{i}=i\cdot\max_{k,u^{k}\geq i}p^{k}$ for each $i$.
Consider now any chain $\mathcal{C}^{k}$. In any schedule $s$, $S^{k}(s)\geq n^{k}p^{k}\geq n^{k}p^{\min}$. And with the proposed offset schedule $\omega$, $S^{k}(\omega)=(n^{k}-1)p^{\max}+p^{k}\leq n^{k}p^{\max}$. So, $\frac{S^{k}(\omega)}{S^{k}(s)}\leq\frac{p^{\max}}{p^{\min}}$.
Now, for some chain $\mathcal{C}^{k}$,
$D_{\max}(\omega)=\left\lceil\frac{S^{k}(\omega)}{T^{k}}\right\rceil-1\leq\frac{S^{k}(\omega)}{T^{k}}\leq\frac{p^{\max}}{p^{\min}}\frac{S^{k}(opt)}{T^{k}}\leq\frac{p^{\max}}{p^{\min}}(D_{\max}(opt)+1)$
If we assume harmonic periods and processing times (making the single resource problem polynomial), Lemma 2 gives a simple polynomial algorithm to compute a good offset schedule for criteria $D_{sum}$ and provides a performance guarantee for criteria $D_{\max}$.
When processing times are arbitrary, we can still apply the offset schedule as long as we are able to find a feasible schedule using, e.g., CP or SMT. First, we find a feasible schedule for the first resource of the line, and then we create schedules for other resources by applying a suitable offset $o_{i}$. Even though the single resource problem remains strongly NP-hard, we empirically show in Section 5.3 that the offset schedule combined with the CP model for the single resource problem works very well.
3.2.2 Periodic Scattering along a Tree
Similar offset schedules can be used on scattering on an out-tree topology, assuming that the first task of each chain is performed on the root resource of the tree and the following tasks use consecutive resources along a path of the tree. This is shown in Fig. 3. Notice that there are actually two root resources, A-B and A-F, which lead to a decomposition of the instance to scattering along a tree (chains $\mathcal{C}^{1},\mathcal{C}^{2}$) and along a line (chain $\mathcal{C}^{3}$). Only the first tasks of the chains $\mathcal{C}^{1},\mathcal{C}^{2}$ share the same resource A-B.
Figure 3: Example of a scattering along a tree instance. Chains $\mathcal{C}^{1},\mathcal{C}^{2}$ originate on link A-B, and their paths diverge afterward. Chain $\mathcal{C}^{3}$ is completely independent since it originates on a different link A-F.
Similarly, as in the previous section, we can duplicate the schedule of the root on the intermediate nodes, removing the chains that are not processed and defining offsets with respect to the schedule of the root resource. Similarly, if resource $j$ is a successor of resource $i$ in the tree, then for a chain $\mathcal{C}^{k}$ that uses this communication link, we should have $o_{j}-o_{i}\geq p_{i}^{k}=p^{k}$. So that the optimal degeneracy of offset schedules based on a given schedule of the root resource is easy to compute. Analogously, this approach can be used for gathering information in the sink of an in-tree.

### v2

@@ 3.1.2 Scattering on a Line of Two Resources with a Single Period
Let us consider a special case of the line support graph with two resources and a scattering hypothesis. So, all chains have their first task on resource 1 and second on resource 2. Assume all chains have the same period $T$. Notice that in this case, the sum of processing times of tasks on any resource is not greater than $T$. Otherwise, the problem would be infeasible. Then, finding a feasible schedule for a single resource becomes trivial. We can prove that as for the two-machine flow-shop problem, minimizing the degeneracy can be done in polynomial time. The same is true for deciding the existence of a periodic schedule with no wait time.
Property 2.
For any instance of a scheduling problem $PD2|T_{i}=T,jit=0,chains|D_{sum}$, in the special case of scattering on a line of two resources:
  • 1.
    Assuming that the sum of processing times on each machine is not greater than period $T$, there always exists a feasible schedule with
    $D_{sum}=D_{\max}=0$ (30)
  • 2.
    It can be decided in polynomial time whether exists a feasible schedule with a no-wait hypothesis (i.e., where $S^{k}=p_{1}^{k}+p_{2}^{k}$ for each chain $\mathcal{C}^{k}$)
Proof.
The non-periodic version of this problem is a two-machine flow-shop. A chain $\mathcal{C}^{k}$ corresponds to a job with two operations $J_{1}^{k},J_{2}^{k}$ performed on the first and second machine, respectively. As shown in [41], a schedule that optimizes the makespan ($C_{max}=\max_{k\in\{1,\ldots,\kappa\}}t_{2}^{k}+p_{2}^{k}$) can be computed in polynomial time using Johnson’s rule. Moreover, we know that it is a permutation schedule: the order of jobs is the same on both machines even though it was not required.
If we shift the schedule of the first machine to the left and of the second machine to the right so that there does not remain idle time between two tasks on the resources and it ends at the optimal makespan, then the schedule can be repeated every $T$ time units without conflicts. Let $t$ be the Johnson’s schedule shifted as indicated above.
We prove that this leads to a zero degeneracy periodic schedule. Indeed, consider a pivot chain $\mathcal{C}^{k^{*}}$, for which $t_{1}^{k^{*}}+p_{1}^{k^{*}}=t_{2}^{k^{*}}$. Such a pivot chain exists. Otherwise, the tasks on the second machine could be scheduled a little earlier, and thus, the makespan would not be optimal. The start-to-end latency for chain $\mathcal{C}^{k^{*}}$ is thus minimal, and so is its degeneracy. Then, for any chain $k$ performed before this $k^{*}$ in the schedule, the start-to-end latency $S^{k}\leq t_{2}^{k^{*}}$ is not greater than the sum of processing times on the first resource, which is not greater than $T$. Hence $D^{k}=0$. Symmetrically, for a chain $l$ performed after chain $k^{*}$, the value of $S^{l}$ should be less than the sum of processing times on the second resource, which is not greater than $T$ so that $D^{l}=0$. This proves that in this schedule, we have $D_{\max}=D_{sum}=0$.
Now consider the no-wait problem where we wish that for each chain $\mathcal{C}^{k}$ the start-to-end latency is $S^{k}=p_{1}^{k}+p_{2}^{k}$. As proved in [42], a no-wait periodic schedule with minimal period $T^{\prime}$ that minimizes the idle time on the second machine can be built using a polynomial subcase of the traveling salesman problem described in [43]. So if the period $T$ is not less than $T^{\prime}$, we can just delay the iterations to get a periodic schedule with period $T$. Otherwise, the problem is not feasible. ∎
Property 2 indicates that minimizing $D_{sum}$ and $D_{\max}$ is polynomial on two machines with a single period. Still, the complexity of minimizing $\sum_{k}S^{k}$ or $\max_{k}S^{k}$ are open problems and might be more difficult if we consider that flow-shop problem on two machines with limited buffers is NP-hard problem (see survey [44]).
3.1.3 Scattering on a Line of Two Resources with Two Periods
As in Section 3.1.2, we assume scattering chains on two resources, but now the chains can have two different harmonic periods. In [7], the authors give a reduction from $3$-PARTITION to a single resource scheduling problem with two different harmonic periods, thus proving that the feasibility of single resource PSP with two periods is strongly NP-hard.
From this result, we can deduce the following:
Property 3.
For scattering chains on two resources with two harmonic periods on a line, finding a schedule with $D_{sum}$ or $D_{\max}$ equal to zero is strongly NP-hard.
Proof.
Consider an instance of the single resource scheduling problem with two harmonic periods. We build a simple reduction to the problem on two resources. We are given a set $T$ of tasks, a task $J^{k}\in T$ has period $T^{k}\in\{\tau(1),\tau(2)\}$, a set of two harmonic periods. Notice that the processing time of $J^{k}$ can be assumed strictly less than its period; otherwise, $J^{k}$ would be the only task scheduled on the resource. For each task $J^{k}\in T$, we define a chain $\mathcal{C}^{k}$ with period $T^{k}$ composed of task $J_{1}^{k}=J^{k}$ on the first resource and a task $J_{2}^{k}$ with processing time $1$ on the second resource. Obviously, if a feasible schedule for the single resource instance exists, it gives a schedule for the two resources of the chains with 00 degeneracy by scheduling the tasks on the second resource as soon as possible with respect to precedence constraint. Conversely, a schedule with 00 degeneracy $D_{sum}$ or $D_{\max}$ implies the existence of a feasible schedule on the first resource. So, scattering problem on two resources with two periods is NP-hard in the strong sense. ∎
As was shown in [7], the single resource problem with processing times and the periods belonging to a harmonic set is polynomial. However, the complexity of building zero-degeneracy schedules for the two-resource scattering on a line where periods and processing times are from one harmonic set remains open.
3.2 Acyclic Instance with Bottleneck Resource and Equal Processing Times
In this section, we assume equal processing times of all tasks of a chain and that the support graph is acyclic and contains a bottleneck resource $b$. We first consider the special case of a line support graph and then extend it to other graph structures.
3.2.1 Bottleneck Resource along a Line
We consider a case when messages are sent in a single direction with a line support graph and a single bottleneck resource and equal processing times of all tasks of a chain. More formally, we consider a support graph that forms a unique directed path from resource $1$ to resource $M$ and a bottleneck resource $b\in\{1,\ldots,M\}$. We consider that all chains $\mathcal{C}^{k}$ have a source resource $u_{k}\leq b$ that performs the first task of the chain ($\pi_{1}^{k}=u_{k}$) and a sink resource $v_{k}\geq b$ that performs the last task of the chain ($\pi^{k}_{n^{k}}=v_{k}$), and if $J_{i}^{k}$ is performed on some resource $m$, with $i<n^{k}$, then $J_{i+1}^{k}$ is performed on resource $m+1$. Notice that scattering corresponds to the case where $b=1$. Let us denote by $p^{k}$ the processing time of every task of chain $\mathcal{C}^{k}$.
Let us assume that we have a schedule $\sigma$ for resource $b$, which contains one task of each chain. We define a special kind of schedules, called offset schedules, that define schedules of other resources according to $\sigma$, with an offset depending only on the resource.
Definition 13 (Offset schedule).
An offset schedule associated with a schedule $\sigma$ on resource $b$ is a schedule built by assuming that for any chain $\mathcal{C}^{k}$ and any resource $m\in\{u_{k},\ldots,v_{k}\},m\neq b$, a $J_{i}^{k}\in\mathcal{J}_{m}$ is scheduled exactly as it is in resource $b$, with an offset $o_{m}$ that only depends on $m$, and can be negative or positive. If $J_{j}^{k}$ is the task of chain $\mathcal{C}^{k}$ performed on resource $b$, the start time of $J_{i}^{k}$ is thus $t_{i}^{k}=\sigma_{j}^{k}+o_{m}$.
Notice that as long as $\sigma$ satisfies the constraint on resource $b$, any offset schedule based on $\sigma$ also satisfies the resource constraints on the other resources. The start-to-end delay of a chain $\mathcal{C}^{k}$ in an offset schedule is $S^{k}=p^{k}+o_{v_{k}}-o_{u_{k}}$. The next property establishes that an offset schedule with an approximation ratio can be built.
Property 4.
Suppose that all chains start before a bottleneck resource $b$ and end after $b$, along a line, assuming equal processing times within each chain. Suppose we are given a feasible schedule $\sigma$ of the tasks on the bottleneck resource $b$. The minimal degeneracy $D_{\max}$ or $D_{sum}$ of an offset schedule $\omega$ associated with $\sigma$ can be computed in polynomial time. In this schedule, we have the following inequality:
$D_{\max}(\omega)\leq\frac{p^{\max}}{p^{\min}}(D_{\max}(opt)+1),$ (31)
where $p^{\min}=\min_{k}p^{k},p^{\max}=\max_{k}p^{k}$ and $D_{\max}(opt)$ is the optimal $D_{\max}$.
Proof.
We first establish the necessary conditions for offsets to give a feasible schedule; then, we prove that we can determine the offsets so that the start-to-end delay of chains is minimal (among offset schedules). Eventually, we show the performance ratio by using a lower bound on the star-to-end latency.
Precedence relations imply that in any feasible offset schedule based on schedule $\sigma$, we have the following inequality:
$\text{If}\ m>b,\quad o_{m}-o_{m-1}\geq\max_{k,v_{k}\geq m}p^{k}$ (32)
Indeed, if $J_{i}^{k}$ is the task of chain $\mathcal{C}^{k}$ performed on resource $m$ ($v_{k}\geq m$), then due to the line support graph $J_{i-1}^{k}$ is performed on resource $m-1$. Hence $t_{i}^{k}\geq t_{i-1}^{k}+p^{k}$, which proves that:
$o_{m}-o_{m-1}\geq p^{k}$ (33)
With similar arguments, we can prove that:
$\text{If}\ m<b,\quad o_{m+1}-o_{m}\geq\max_{k,u_{k}\leq m}p^{k}$ (34)
Indeed, if $J_{i}^{k}$ is performed on resource $m$ ($u_{k}\leq m$), $J_{i+1}^{k}$ is performed on resource $m+1$, so that $t_{i+1}^{k}\geq t_{i}^{k}+p^{k}$. So:
$o_{m+1}-o_{m}\geq p^{k}$ (35)
We can now build a feasible offset schedule $\omega$ that minimizes the start-to-end latency of each chain (among offsets schedules) by setting $o_{b}\coloneqq 0$ and if $m>b$ then $o_{m}\coloneqq o_{m-1}+\max_{k,v^{k}\geq m}p^{k}$, whereas if $m<b$, then $o_{m}\coloneqq o_{m+1}-\max_{k,u^{k}\leq m}p^{k}$. This can be computed in time complexity $\mathcal{O}(\mu log(\mu))$ by sorting the processing times and, for at most $M$ steps (from $b$ to $M$ and from $b-1$ to $1$), scanning the list to find the maximum processing time concerned by resource $m$.
Then, for a chain $\mathcal{C}^{k}$ it holds:
$o_{v_{k}}-o_{b}\leq(v_{k}-b)p^{max},$ (36)
whereas:
$o_{b}-o_{u_{k}}\leq(b-u_{k})p^{max}.$ (37)
The start-to-end latency of $\mathcal{C}^{k}$ in this schedule satisfies:
$S^{k}(\omega)\leq o_{v_{k}}+p^{k}-o_{u_{k}}\leq(v_{k}-u_{k}+1)p^{max}$ (38)
To conclude, we observe that the start-to-end latency of chain $\mathcal{C}^{k}$ in any schedule is not less than $(v_{k}-u_{k}+1)p^{min}$. So, for any schedule $s$, $\frac{S^{k}(\omega)}{S^{k}(s)}\leq\frac{p^{\max}}{p^{\min}}$.
Now, for some chain $\mathcal{C}^{k}$:
$D_{\max}(\omega)=\left\lceil\frac{S^{k}(\omega)}{T^{k}}\right\rceil-1\leq\frac{S^{k}(\omega)}{T^{k}}\leq\frac{p^{\max}}{p^{\min}}\frac{S^{k}(opt)}{T^{k}}\leq\frac{p^{\max}}{p^{\min}}(D_{\max}(opt)+1)$ (39)
In the case of processing times and periods belonging to a harmonic set, a schedule $\sigma$ of the bottleneck resource can be found in polynomial time. Then Property 4 gives a polynomial approximation algorithm in this case, assuming an acyclic line support graph with a bottleneck resource, equal processing times within chains, and harmonic processing times and periods.
When processing times are not within a harmonic set with the periods, we can still build an offset schedule as long as we are able to find a feasible schedule using, e.g., CP. First, we find a feasible schedule for the first resource of the line, and then we create schedules for other resources by applying a suitable offset. Even though the single resource problem remains strongly NP-hard, we empirically show in Section 5.3 that the offset schedule combined with the CP model for the single resource problem works very well.
3.2.2 Bottleneck Resource on Any Acyclic Graph
We now consider that the support graph $G=(V,E)$ does not contain any directed cycle and that there is a bottleneck resource $b$. We show that the previous result can be extended to this case. As previously, we assume a feasible schedule $\sigma$ on bottleneck resource $b$. Two consecutive tasks of chain $\mathcal{C}^{k}$, $J_{i}^{k},J_{i+1}^{k}$ with $\pi_{i}^{k}=m$, $\pi_{i+1}^{k}=m^{\prime}$ then satisfy that $(m,m^{\prime})\in E$ is an arc of the support graph. We say that the chain passes through arc $(m,m^{\prime})$. For an arc $a=(m,m^{\prime})\in E$ of the support graph, let us define by $C(a)$ the set of chains $\mathcal{C}^{k}$ that pass through arc $a$.
In Fig. 4, we present an example of an instance with an acyclic support graph and a bottleneck resource 3. Notice that the support graph contains an undirected cycle (resources 3, 4, 5, 7, 8). This does not invalidate the results of this section since there is still no directed cycle.
A crucial distinction between the network topology and the support graph can be observed in Fig. 4 by focusing only on network nodes A, B, and C (and removing nodes D-H). In that case, even though the messages share the sink network node C, the support graph, consisting only of resources 1 and 2, does not contain any edges (and no bottleneck resource). Thus, such PSP decomposes into scheduling chains $\mathcal{C}^{1},\mathcal{C}^{2}$ on resource 1 and independently scheduling chain $\mathcal{C}^{3}$ on resource 2.
Figure 4: Example of acyclic PSP with a bottleneck resource. Underlying network topology (blue) induces an appropriate support graph (dashed) given the three chains of tasks (green) starting at links A-C and B-C, passing through bottleneck link C-D.
In the remainder of the subsection, we do not make other specific assumptions on the support graph’s structure. However, we will show that the performance of offset schedules can be measured with respect to a parameter $\eta$.
Definition 14 (Ratio between path lengths).
Parameter $\eta$ measures the maximum ratio between the length of two different paths between two nodes in the support graph $G$. More formally, if $P$ is a directed path of the support graph from node $m$ to node $m^{\prime}$, then its length $\lambda(P)$ is the number of nodes of the path, and $\mathcal{P}(m,m^{\prime})$ denotes the set of paths from $m$ to $m^{\prime}$ in $G$. We also denote by $m\rightarrow m^{\prime}$ the fact that $\mathcal{P}(m,m^{\prime})\neq\emptyset$. Then
$\eta=\max_{m,m^{\prime}\in V,m\rightarrow m^{\prime}}\frac{\max_{P\in\mathcal{P}(m,m^{\prime})}\lambda(P)}{\min_{P\in\mathcal{P}(m,m^{\prime})}\lambda(P)}$ (40)
Notice that in several important special cases, for example, line and tree support graphs (or when any two paths from $m$ to $m^{\prime}$ have the same length), we have $\eta=1$. This does not hold for Fig. 4, where there are multiple paths between resources 3 and 8. As previously mentioned, an offset schedule will define the schedule of resource $m$ based upon $\sigma$ with an offset $o_{m}$, assuming offset of the bottleneck resource $o_{b}=0$. We can now prove the following property of offsets:
@@ 3.2.2 Bottleneck Resource on Any Acyclic Graph
For any arc $a=(m,m^{\prime})\in E$ of the support graph, in any offset schedule:
$o_{m^{\prime}}-o_{m}\geq\max_{k,\mathcal{C}^{k}\in C(a)}p^{k}$ (41)
@@
Suppose that a chain $\mathcal{C}^{k}\in C(a)$. Then $J_{i}^{k},J_{i+1}^{k}$ are consecutive in the chain with $\pi_{i}^{k}=m,\pi_{i+1}^{k}=m^{\prime}$. Suppose that $J_{j}^{k}$ is the task of chain $\mathcal{C}^{k}$ performed on resource $b$. The precedence constraint between $J_{i}^{k},J_{i+1}^{k}$ implies that:
$o_{m}+\sigma_{j}^{k}+p^{k}\leq o_{m^{\prime}}+\sigma_{j}^{k}.$ (42)
So we can now consider the support graph $G$ with its arc valuation $v$: for each arc $a=(m,m^{\prime})\in E$ $v(a)=\max_{k,\mathcal{C}^{k}\in C(a)}p^{k}$. The valuation $v(P)$ of a path $P$ of $G$ is defined as the sum of the valuation of its arcs. From this valuation we define the following offsets $\hat{o}_{m}$:
Definition 15 (Maximal path offsets $\hat{o}$).
If there is a path from $b$ to $m$ in G, then we define :
$\hat{o}_{m}=\max\{v(P)|P\ path\ from\ b\ to\ m\}$ (43)
Otherwise:
$\hat{o}_{m}=\max\{-v(P)|P\ path\ from\ m\ to\ b\}$ (44)
Theorem 2.
In the case of acyclic instance with a bottleneck resource $b$, and given a schedule $\sigma$ of resource $b$, the offset schedule $\omega$ defined upon maximal path offsets $\hat{o}_{m}$ can be computed in polynomial time from $\sigma$ and has the property that
$D_{\max}(\omega)\leq\eta\cdot\frac{p^{\max}}{p^{\min}}(D_{\max}(opt)+1),$ (45)
where $p^{\min}=\min_{k}p^{k},p^{\max}=\max_{k}p^{k}$ and $D_{\max}(opt)$ is the optimal $D_{\max}$.
@@
We first prove that any offset schedule satisfies that the offset difference along a directed path induced by a chain $\mathcal{C}^{k}$ gives a potential inequality. Then, we prove that offsets $\hat{o}$ satisfy these inequalities and can be computed in polynomial time. We then measure the start-to-end latency of a chain in schedule $\omega$ and give a lower bound on the start-to-end latency of a chain that induces our approximation ratio for the degeneracy.
Let us consider a chain $\mathcal{C}^{k}$ starting on resource $m=\pi_{1}^{k}$ and ending at resource $m^{\prime}=\pi^{k}_{n^{k}}$. Since $b$ is a bottleneck resource, $\mathcal{C}^{k}$ induces a directed path $P$ from resource $m$ to $b$ in $G$ and a path $P^{\prime}$ from $b$ to $m^{\prime}$. By Lemma 41, summing the inequalities on the paths implies that $o_{m^{\prime}}-o_{b}\geq v(P^{\prime})$ whereas $o_{b}-o_{m}\geq v(P)$. By definition, these inequalities are satisfied by $\hat{o}$ values. Paths with maximum value can be computed in polynomial time ($\mathcal{O}(M^{2})$ in the acyclic graph $G$). Now let us consider the start-to-end latency of chain $\mathcal{C}^{k}$ for schedule $\omega$. As in Property 4 we get $S^{k}(\omega)=\hat{o}_{m^{\prime}}+p^{k}-\hat{o}_{m}$. Now observe that the maximum value of a path $P$ from $m$ to $m^{\prime}$ plus the processing time $p^{k}$ is not greater than $\max_{Q\in\mathcal{P}(m,m^{\prime})}\lambda(Q)p^{\max}$, whereas the optimal start-to-end latency of chain $\mathcal{C}^{k}$ is not less than $(n^{k}+1)p^{min}\geq\min_{Q\in\mathcal{P}(m,m^{\prime})}\lambda(Q)p^{\min}$. So we get
$S^{k}(\omega)\leq\eta\cdot\frac{p^{\max}}{p^{\min}}S^{k}(opt)$ (46)
We can then use the same arguments as in the proof of Property 4 to conclude. ∎
Corollary 2.
In the case of a tree support graph and a bottleneck resource $b$, and processing times and periods belonging to a harmonic set, a schedule $\omega$ with the following approximation ratio can be computed in polynomial time.
@@
We just have to build, using Korst’s algorithm [7], a schedule $\sigma$ for the bottleneck resource and then compute the optimal offset schedule associated with $\sigma$ along with Theorem 2. ∎